OpenStudy (anonymous):

what is the equation of a circle with the center at (-3,1) and through the point (2, 13)

OpenStudy (mathstudent55):

Remember the standard equation of a circle from the previous problem?

OpenStudy (mathstudent55):

\( (x - h)^2 + (y - k)^2 = r^2 \)

OpenStudy (anonymous):

ok so h= -3 and k=1

OpenStudy (mathstudent55):

Right. Let's put that in the equation: \((x + 3)^2 + (y - 1)^2 = r^2\) Ok?

OpenStudy (mathstudent55):

Now take the given point, (2, 13), and replace x with 2 and y with 13. Then you can find r. Then rewrite the equation using the r value you find.

OpenStudy (anonymous):

so (2+3)^2 + (13-1)^2 =11^2 right?

OpenStudy (mathstudent55):

How do you know r = 11? That's what you are solving for.

OpenStudy (anonymous):

(2+3)^2 + (13-1)^2= r^2

OpenStudy (mathstudent55):

\( (2+3)^2 + (13-1)^2 =r^2\) Solve for r.

OpenStudy (mathstudent55):

Correct. Calculate the number on the left side and find r.

OpenStudy (anonymous):

ok so (2+3)^2=11 or do i do it 2+3 they square it making it 25

OpenStudy (mathstudent55):

Do what is in parentheses first. Then square.

OpenStudy (anonymous):

so it would be r= 25^2

OpenStudy (mathstudent55):

What happened to (13 - 1)^2?

OpenStudy (mathstudent55):

\( (2+3)^2 + (13-1)^2 =r^2\) \(5^2 + 12^2 = r^2 \) \(25 + 144 = r^2\) \(169 = r^2\) r = ?

OpenStudy (anonymous):

13

OpenStudy (mathstudent55):

Good. Now that we know r, we can write the final solution: The equation of the circle is \((x+3)^2 + (y-1)^2 =169\)